A woman on the earth firecracker explode 10.0m in front of her when her clock reaches 5.0 seconds an astronaut in rocketship who passes the woman on earh at t=0 at speed of 0.40c. Find what coordinate for this event?


1
Expert's answer
2021-02-18T18:56:05-0500

Explanations & Calculations


  • Since the rocket is flying with relativistic speed, the time taken for the firecracker to explode is detected differently by the women and the astronaut.
  • This phenomenon is described by the equation,

"\qquad\qquad
\begin{aligned}
\small T&= \small \frac{1}{\sqrt{1-\frac{v^2}{c^2}}}\cdot\tau
\end{aligned}"

where T represents the time seen by the earth bounder & "\tau" represents time seen by the one in the reference frame moving at relativistic speed.

  • Therefore,

"\qquad\qquad
\begin{aligned}
\small 5\,s &= \small \frac{1}{\sqrt{1-0.4^2}}\cdot\tau\\
\small \tau &= \small 5\times 0.916515139\\
&= \small 4.583\,s
\end{aligned}"

  • The astronaut sees the cracker explode after 4.583 s after he passes the women